You have a 12 V battery connected to a single straight nichrome wire in your heater. if the wire is a millimeter thick and a meter long and you place a devise that measures magnetic stregth at a one centimeter distance away from the wire, what will be the reading on this devise?
1
Expert's answer
2021-09-14T15:07:32-0400

Nichrome wire has resistivity

"ρ = 1 \times 10^{-6} \; ohm \cdot m \\

l = 1 \;m \\

A = \pi (\frac{d}{2})^2 \\

A = \pi (\frac{1 \times 10^{-3}}{2})^2 \\

A = 7.85 \times 10^{-7} \;m^2"

Resistance of this wire will be

"R = \frac{ρl}{A} \\

R= \frac{1 \times 10^{-6} \times 1}{7.85 \times 10^{-7}} \\

R = 1.27 \; ohm"

Current I in this qire will be

"I = \frac{V}{ R} \\

I = \frac{12}{1.27} \\

I = 9.44 \;A"

Magnetic field at distance R from a wire is given by

"B = \frac{\mu_0 I}{2 \pi R} \\

B = \frac{4 \pi \times 10^{-7} \times 9.44}{2 \pi \times 1. \times 10^{-2} } \\

B = 1.88 \times 10^{-4} \;T"